OpenStudy (anonymous):

\[y=\frac{ 1 }{ 2}x-1\] Complete the ordered pairs (2, ) (0, ) Help please !!

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started

jimthompson5910 (jim_thompson5910):

plug in x = 2 \[\large y=\frac{ 1 }{ 2}x-1\] \[\large y=\frac{ 1 }{ 2}(2)-1\] \[\large y=1-1\] \[\large y=0\] Therefore, the ordered pair (2,0) is on the line \[y=\frac{ 1 }{ 2}x-1\]

OpenStudy (anonymous):

ok im going to attempt the other one

OpenStudy (anonymous):

-1 ?

jimthompson5910 (jim_thompson5910):

you got it, so (0,-1)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Use a proportion to solve this problem. Jim drove 250 miles in 5 hours. If he can keep the same pace, how long will it take him to drive 1,100 miles ?

OpenStudy (anonymous):

can you help me with this